### v1

# Q2E: Query-to-Event Decomposition for Zero-Shot Multilingual Text-to-Video Retrieval

Abstract
Recent approaches have shown impressive proficiency in extracting and leveraging parametric knowledge from Large-Language Models (LLMs) and Vision-Language Models (VLMs). In this work, we consider how we can improve the identification and retrieval of videos related to complex real-world events by automatically extracting latent parametric knowledge about those events. We present Q2E: a Query-to-Event decomposition method for zero-shot multilingual text-to-video retrieval, adaptable across datasets, domains, LLMs, or VLMs. Our approach demonstrates that we can enhance the understanding of otherwise overly simplified human queries by decomposing the query using the knowledge embedded in LLMs and VLMs. We additionally show how to apply our approach to both visual and speech-based inputs. To combine this varied multimodal knowledge, we adopt entropy-based fusion scoring for zero-shot fusion. Through evaluations on two diverse datasets and multiple retrieval metrics, we demonstrate that Q2E outperforms several state-of-the-art baselines. Our evaluation also shows that integrating audio information can significantly improve text-to-video retrieval. We have released code and data for future research. [^1] https://github.com/dipta007/Q2E
Q2E: Query-to-Event Decomposition for Zero-Shot Multilingual Text-to-Video Retrieval
Shubhashis Roy Dipta, Francis Ferraro Department of Computer Science and Electrical Engineering University of Maryland Baltimore County Baltimore, MD 21250 USA {sroydip1,ferraro}@umbc.edu
2 Introduction
Figure 1: In Q2E, we extract prequel, current, and sequel events, along with an audio transcript and video description to enrich the query context and enhance video retrieval. This information matches against both visual, textual, and speech-based descriptions (matching phrases are highlighted in the same color) enabling the retrieval of the correct videos while effectively filtering out a visually similar but non-relevant video.
Widely used video retrieval websites can give the impression that video retrieval is a solved problem: a user types in a query, and gets back a nicely curated list of relevant videos. However, such approaches can leverage rich metadata, or subsequent human-creator curation (such as titles, or search-optimized descriptions), that go beyond the information conveyed directly in that video. For example, as shown in Fig. 1, there are many different videos of fires, but if a user provides a basic query such as “2025 LA fire” in order to find videos providing information about that specific fire, how can a system satisfy this request? In this paper, we present an approach we call Q2E (query-to-event decomposition) that leverages the embedded knowledge that LLMs have about complex events, and use this knowledge to automatically decompose and elaborate a user’s query, while simultaneously using this knowledge to identify salient visual and audio-based features from the video.
Recent work (Liu et al., 2024b) has found that users typically input basic queries, expecting the system to capture “fine-grained semantic concepts” and understand complex relationships between videos and text. Our proposed approach decomposes queries into multiple fine-grained events, providing critical elaboration of the implicit information requests of queries. Additionally, most current text-to-video retrieval systems are trained and evaluated on widely used datasets, such as MSR-VTT (Xu et al., 2016) and MSVD (Chen and Dolan, 2011), which feature generic, high-level queries like “a person is explaining something” (a random query from MSR-VTT). These generic queries restrict text-to-video retrieval systems’ learning capacity, preventing them from effectively retrieving specific videos based on precise user queries, i.e., “2025 LA Fire”. Finally, most text-to-video retrieval datasets focus on English language videos, whereas videos transcend linguistic boundaries, necessitating retrieval methods adaptable to other languages.
Motivated by these challenges, we explore how to use query expansion and video+audio features to retrieve highly specific videos across different languages. Furthermore, we explore how the extensive knowledge embedded in Large-Language Models (LLMs) and Vision-Language Models (VLMs) can be effectively transferred to text-to-video retrieval. To achieve this, we propose a novel zero-shot text-to-video retrieval method that integrates the pre-existing knowledge of LLMs and VLMs to enhance text-to-video retrieval.
Underlying our approach are three key insights: (1) Transferring knowledge from LLMs through query decomposition enhances the understanding of coarse-grained human queries. This provides richer context and enables the retrieval of prequel and sequel videos that would otherwise be overlooked. (2) While VLM-based captioning and large speech ASR models can generally describe an image/video frame and transcribe a portion of speech, the outputs can be repetitive, too low-level, missing broader themes that occur across frames. However, we can use an LLM to remove redundancy, and automatically summarize these lower-level descriptions for empirical effectiveness. (3) In our approach, determining whether an individual video is relevant requires aggregating across multiple forms of similarity and relevancy judgments, each of varying strength. Entropy-based rank fusion provides an intuitive way of performing this aggregation and ranking, while surpassing both simple fusion methods (mean, max) and more complex approaches like Reciprocal Rank Fusion. Using these insights, Q2E achieves strong performance gains on two varied and challenging datasets.
In summary, our contributions are as follows:
  • We propose a novel framework that leverages the embedded world knowledge of LLMs to enrich human queries with prequel, current, and sequel-based events.
  • We use large video, speech, and language models to capture critical event- and query-relevant information communicated across those multiple modalities, how to assess that information for query relevance, and how to combine those assessments via entropy-based rank fusion.
  • Through extensive ablation studies, we have shown the importance of each component of our method in the final performance.
  • Our zero-shot method improved performance on the widely used text-to-video retrieval dataset, MSR-VTT (Xu et al., 2016) (+4 NDCG points above our baseline), as well as on the more enriched, multi-lingual event-focused retrieval dataset, MultiVENT (Sanders et al., 2023) (+8 NDCG points).
3 Related Work
With the rapid advancements in LLMs (Devlin et al., 2019; Radford et al., 2019; Touvron et al., 2023; Dubey et al., 2024; Gunasekar et al., 2023; Liu et al., 2024a; Jiang et al., 2024; Chu et al., 2024), methodologies such as zero-shot, few-shot, and in-context learning have gained substantial traction due to their reduced computational requirements and enhanced time efficiency. These have influenced approaches for event causal reasoning, multimodal language models, and retrieval.
3.1 Event Causal Reasoning
Reasoning based on event causality plays a crucial role across various domains. Sun et al. (2024) demonstrated that event-causal graphs, even generated by LLMs, can enhance story understanding and better align with human evaluations. Similarly, in Q2E, we leverage causation from human queries to construct prequel, current, and sequel events, enabling a deeper understanding of user intent. While Chan et al. (2024) utilized LLMs to infer temporal, causal, and discourse relationships between two events, our approach applies the same temporal (prequel or sequel) and causal (current) reasoning to extract relevant events directly from queries.
3.2 Vision Language Model
Vision-Language Models (VLMs) have been used significantly in text-to-video retrieval tasks by leveraging large-scale pre-trained models trained on image-text pairs (Ibrahimi et al., 2023). These models have demonstrated impressive zero-shot capabilities and transfer learning potential for video-related tasks (Zhao et al., 2024). Recent advancements have focused on efficiently adapting image-based VLMs to the video domain, addressing temporal dynamics and computational costs (Nishimura et al., 2024). Some approaches propose novel architectures like sparse-and-correlated adapters (Cao et al., 2024) to enhance retrieval performance while maintaining efficiency. Additionally, researchers have explored incorporating audio information (Ibrahimi et al., 2023) and using synthesized instructional data (Zhao et al., 2024) to improve video-language models further.
3.3 Zero-Shot Text-to-Video Retrieval
Though much has been done in the domain of fine-tuned text-to-video retrieval (Wang et al., 2025; Cicchetti et al., 2024; Chen et al., 2023; Xu et al., 2023), there is a gap in zero-shot text-to-video retrieval performance. Fine-tuning a model on a task-specific dataset can lead to impressive performance, but it is not always suitable due to high computing costs and time constraints. While zero-shot text-to-video retrieval has been previously studied (Wang et al., 2025; Cicchetti et al., 2024; Chen et al., 2023; Xu et al., 2023) results suggest a sizable gap between that and fine-tuned retrieval.
MSR-VTT (Xu et al., 2016) is one of the well-known benchmarks. However, the queries in the dataset can tend to be more generic, e.g., “a tornado and a thunderstorm are looming in the background,” rather than targeting rich information about a specific event, like “2025 LA Fire” from Fig. 1. In contrast, MultiVENT (Sanders et al., 2023) defines a multilingual video retrieval task where the queries refer to specific events.
4 Q2E: Query-to-Event Decomposition
Figure 2: The complete method of our text-to-video retrieval system. The top blue box represents the query decomposition module of our method, while the bottom green box illustrates the video captioning and audio transcription module. The left orange box represents our multi-layer ASR module. The right purple box fuses the ranks calculated from query-video, query-descriptions, and event-descriptions (event = prequel + current + sequel, description = multimodal description). Text-to-text similarity scores were computed using ColBERT (Khattab and Zaharia, 2020), and text-to-video similarity scores were computed using MultiCLIP (Delitzas et al., 2023). In the non-ASR variant of our method, the components within the orange box and orange dotted lines are excluded.
Modern dense retrieval approaches commonly compute an embedding-based similarity (e.g., cosine) between a query and each candidate; those scores rank candidates, and the top-scoring candidates are returned. For example, in video retrieval, one could embed the user’s query using an appropriate LLM, embed the video using an appropriate video embedder, and then compute the cosine similarity between those embeddings Sanders et al. (2023). This “query vs. video” similarity score is used as a proxy for how relevant to video is to the query.
Our approach (shown in Fig. 2) builds upon this general embedding similarity computation, though to better capture the diverse and nuanced multimodal information conveyed in videos and human queries—including the event decompositions Q2E produces—we must extend those similarity computations. In order to capture the rich event information we get from our decomposition (§ 4.1), a single embedding representing the video is not sufficient. We need to additionally get multimodal descriptions of the video (§ 4.2 and 4.3); these are text-based descriptions of the video and audio in the video, which can be more readily evaluated against the query. Using these components, Q2E computes five types of scores $S_{i,v}$ for each video $v$: (a) a query vs. video score, as described above; (b) an event decomposition vs. multimodal description score for each of the prequel, current, and sequel decompositions; and (c) a query vs. multimodal description score. While (a) helps to identify coarse-grained features such as “fire” or “flood”, both (b) and (c) provide fine-grained contextual details. By aggregating these components using entropy-based fusion (§ 4.4), they complement each other to improve the retrieval of specific video events.
4.1 Decomposing Query Events
Like any traditional text-video retrieval system, a user provides a natural language query, which can be a couple of words to 10+ words. In this method, we have used an LLM to decompose it into three sub-events: (1) Prequel, (2) Current, and (3) Sequel. Based on initial experiments, and to manage computation, we only considered five each of prequel, current, and sequel events. We provide our template prompts for each in § A.4.
Prequel
“Prequel” primarily refers to events that can lead to the current event described in the query. An event can have multiple prequels; for example, a wildfire can have strong wind, high-temperature alert, and dry lightning as prequels. Prequel events may enable the event in the query, but do not necessarily cause it.
Current
“Current” mainly refers to the simple decomposed events that occur during the event described in the query. This method decomposes the queries into sub-events likely to be observed. E.g., “2025 LA Fire” can be decomposed into “Building on Fire”, or “Smoke everywhere.”
Sequel
“Sequel” events can result from the event described in the query. Like the prequel, one event can lead to multiple sequel events.
Decomposition Refinement
We refine these event decompositions with temporal, spatial, and primary event data. For this, we have used the same LLM to extract the temporal (2025), spatial (LA), and primary event (Fire) information. Although this rich information can be combined templatically, e.g., “P/S/C (2025, LA, Fire),” during initial development experiments, we found that refining it to read as a more natural, human-provided query markedly improved performance. We produced this refined query automatically using an LLM; we show this in § A.4.3, and provide all responses as part of our anonymized data release. [^2] https://github.com/dipta007/Q2E
4.2 Video Decomposition
In practice, we found that embeddings of videos / video frames were not sufficient to capture the varied actions described by our event decomposition through basic similarity computations. While more advanced similarity approaches, such as cross-encoders, may be feasible, they present computational challenges (requiring running the cross-encoder on every video/query pair). To address this, we instead considered how to extract text descriptions (captions) from the videos themselves, with the aim that these descriptions would be sufficiently similar to our query event decomposition. We found two approaches to be effective for capturing complementary local and global elements across the video: (1) generating contextualized descriptions per selected frame, and (2) generating holistic video descriptions.
Contextualized Frame Descriptions
We sampled K=16 frames from the videos (determined by trading off performance vs. computation—see ablation in § A.1.2). We have tried uniform sampling and a scene-change-based sampling method  [^3] https://www.scenedetect.com/ for frame sampling, and found that uniform sampling, despite its simplicity, works better than scene-change-based sampling (ablation in § A.1.3). As the video contains temporal information and by design, video frames should be connected semantically, so rather than captioning each frame in isolation, we used a sliding window (window size=1) approach and provided the VLM model with the previous frame’s caption (when available). Conditioning on previous frame’s captions helped contextualize subsequent captions. Per video, we have K contextualized frame captions.
Video Descriptions
We found that current VLM models can perform dense captioning but miss the “bigger picture.” For instance, a VLM might describe a video of a wildfire with people running and fire everywhere but fail to recognize the higher-level conclusion of “wildfire.” To address this limitation, we provided an LLM model with all 16-frame descriptions from above and asked it to summarize them into a single video-dense caption, preserving the order of the frame captions.
4.3 ASR Augmented Video Captions
Speech is a critical component of many videos, especially when visual elements alone may not be sufficient enough to convey the full context. Despite using a multilingual ASR model (Radford et al., 2023), the multilingual videos posed a unique challenge to get consistent and accurate ASR transcriptions. To address this, we propose a multi-layer translation pipeline (Fig. 2, left orange box). This pipeline leverages three models: (1) a multilingual ASR model, whisper-v3 (Radford et al., 2023), which transcribes audio into both the original language and its English translation; 2) a translation model, NLLB (Costa-jussà et al., 2022) to generate English transcriptions by translating the original transcript; and (3) an LLM, Llama-70B (Dubey et al., 2024), which refines the English translations produced by the previous two models. Early experiments showed all three were critical for sufficient transcription and translation quality.
4.4 Scoring and Score Fusion
Dense Video Embeddings
We utilized the original query and video to compute a similarity score. The MultiCLIP (Delitzas et al., 2023) model was selected due to its superior performance with multilingual videos compared to the CLIP (Radford et al., 2021) model. Although MultiCLIP is traditionally designed for images, we adapted it for video by following the approach of Sanders et al. (2023): uniformly sampling 16 frames and averaging their embeddings to generate the final video embedding. Cosine similarity was applied to compute the similarity score, which was then scaled to the 0–100 range to ensure comparability with other scores: $S_{i,v}|Q=100.0\times\operatorname{multiCLIP}(Q,V),$ where $S_{i,v}|Q$ denotes the score vector for a given query $Q$, and $\operatorname{multiCLIP}(x,y)$ returns the similarity score between x and y.
[…]
MultiVENT
MultiCLIP 9.83 44.32 70.82 88.80 80.77 65.25 0.92 75.34 86.33 22.12 6
MultiCLIP + ASR 10.08 46.64 74.70 90.73 84.32 68.57 0.93 78.76 88.87 22.66 6
MultiCLIP + Q2E 10.24 46.71 75.76 92.28 85.25 69.73 0.95 80.04 89.42 21.13 6
MultiCLIP + Q2E + ASR 10.32 49.00 79.60 93.05 88.73 73.09 0.95 83.24 91.20 16.44 6
[…]
InternVideo2-1B + ASR 9.68 40.48 62.43 87.64 73.59 57.34 0.91 68.30 83.84 109.70 7
InternVideo2-1B + Q2E 9.54 40.88 63.40 87.26 75.21 58.73 0.92 69.15 83.96 53.84 7
InternVideo2-1B + Q2E + ASR 10.24 44.94 70.79 92.28 81.85 65.14 0.95 76.10 88.09 42.81 6
MSR-VTT-1kA
MultiCLIP 43.52 69.05 76.88 43.62 13.85 7.71 0.54 59.72 54.27 20.29 2
MultiCLIP + ASR 45.43 70.75 77.19 45.53 14.19 7.74 0.56 61.22 56.11 31.31 2
MultiCLIP + Q2E 44.52 71.26 79.40 44.62 14.29 7.96 0.56 61.51 55.84 17.91 2
MultiCLIP + Q2E + ASR 46.23 73.37 81.71 46.33 14.71 8.19 0.58 63.59 57.83 18.73 2
[…]
InternVideo2-1B + ASR 53.17 73.27 81.61 53.27 14.69 8.18 0.62 67.08 62.49 23.41 1
InternVideo2-1B + Q2E 53.47 73.57 82.11 53.57 14.75 8.23 0.62 67.16 62.47 16.73 1
InternVideo2-1B + Q2E + ASR 56.28 76.58 83.72 56.38 15.36 8.39 0.65 69.53 65.06 15.85 1
Table 1: Performance comparison on MultiVENT and the MSRVTT “1k-A” split. We use MultiCLIP (per Sanders et al. (2023)) and InternVideo2 (the best-performing method on the leaderboard) as our baseline encoders. Our method (Q2E), even without ASR, improves NDCG by 5–19 points on the MultiVENT dataset and 1-2 points on the MSRVTT dataset. Adding ASR further enhances performance, with an additional 3–7 points on MultiVENT and 2 points on MSRVTT, resulting in a total improvement of approximately 8–26 points on MultiVENT and 3–4 points on MSRVTT, across two different encoders.
Query vs. Multimodal Descriptions
In this approach, we used the original query but replaced the video content with the Multimodal Description Set, yielding a text-to-text similarity score instead of text-to-video. Specifically, we compute the maximum similarity between the query and all multimodal descriptions. Formally:
$S_{i,v}|Q=\max_{c=1:C}Sim(Q,Cap(v)_{c}),$ (1)
where $C$ is the number of descriptions associated with the video, and $Cap(v)$ returns the multimodal description of the video $v$. $Sim(x,y)$ computes the ColBERT444 We experimented with cosine-based SBERT (Reimers and Gurevych, 2019) and found that ColBERT outperforms SBERT in both accuracy and speed. Notably, in our case, the captions generated by LLMs may have some noise. ColBERT mitigates the impact of this noise by leveraging maximum aggregation at the token level, rather than compressing information into a single vector. This approach effectively reduces or ignores noisy elements in the multimodal descriptions. Additionally, prior studies (Warner et al., 2024) have shown that ColBERT’s approach minimizes information loss introduced by the averaging process used in SBERT, further enhancing its performance. similarity score between $x$ and $y$.
Event Decomposition vs. Multimodal Descriptions
We use the refined prequel/current/sequel events from the query and the multimodal description set to compute a many-to-many similarity score. Specifically, we compute the maximum similarity score across all descriptions for each event and subsequently take the maximum similarity score among all the events as the final score. This approach reduces the negative impact of a single noisy decomposition or caption by leveraging the global maximum instead of averaging, which is more vulnerable to noisy data. Formally:
$S_{i,v}|Q=\max_{l=1:L}\max_{c=1:C}Sim(E_{i}(Q)_{l},Cap(v)_{c}),$ (2)
where $L$ is the number of events (prequel / current / sequel) in the query, and $E_{i}()$ returns the respective event decomposition (prequel / current / sequel).
4.4.1 Score Fusion
As Q2E is zero-shot, fusing all the scores presents a challenge. Naively averaging all scores would assume equal importance for each score, which does not reflect real-world scenarios. Different scores contribute differently to the final result depending on the query, e.g., for videos relating to earthquakes, the Sequel vs. Descriptions score carries significantly more weight than the Prequel vs. Descriptions score, as there is no visual indication of an earthquake before it occurs. Yin and Jiang (2024) demonstrated that entropy-based methods outperform others in zero-shot rank fusion tasks. In this entropy ranking, disparate similarity scores for an individual video are normalized to a softmax distribution, where low entropy indicates high confidence, while high inverse entropy corresponds to high confidence. Building on this insight, we use inverse entropy rank fusion to combine the scores.
Specifically, we convert each of the five previously computed similarity scores $S_{i}$ to a weighted distribution $P_{i}=\mathrm{softmax}(S_{i})$, and compute the entropy $H(P_{i})$ of this distribution. This $P_{i}$ provides an initial likelihood weighting across the $V$ videos, according to the $i$th scoring criteria. We rescale $P_{i}$ by its inverse entropy, computing a final score $\hat{S}$ by summing across all rescaled $P_{i}$:
$\hat{S}=\sum_{i}^{5}\frac{1}{H(P_{i})}*P_{i}.$ (3)
By applying Eq. 3, we effectively assign greater weights to scores with lower entropy, emphasizing more confident predictions. This $\hat{S}$ contains an aggregate score for each video; $\hat{S}$ can be sorted to return the ranked list of videos for the query.
5 Results
5.1 Datasets and Implementation
We experimented with the event-centric MultiVENT dataset (Sanders et al., 2023) and the widely-used MSR-VTT dataset (Xu et al., 2016). MultiVENT provides enriched, event-specific human queries, while MSR-VTT contains 10,000 video clips across 20 categories, with each clip annotated with 20 English sentences. We used the standard “1k-A” test split of 1,000 videos for computational constraint, as used in prior works (Jiang et al., 2022; Guan et al., 2023; Jiang et al., 2023a). Dataset details are in Table 13. Full implementation details can be found in § A.3, though briefly, we used Llama-3 as the LLM, InternVL-2.5 as the VLM, and Whisper-large-v3 as the multilingual ASR models. Additionally, we used ColBERT and MultiCLIP for similarity score calculations.
5.2 Zero-Shot Text-to-Video Retrieval
The comparison of our method with the baseline for MultiVENT and MSR-VTT-1kA is shown on the Table 1. We use the standard retrieval metrics: recall/precision at rank K, mean reciprocal rank (MRR), mean average precision (MAP), normalized discounted cumulative gain (NDCG), and mean and median rank (MnR, MdR).
Overall, combined with ASR, Q2E gives the best score in all the metrics across two different datasets and two different text-to-video encoders. But even without audio information, we outperformed the baseline in all metrics. Q2E is also generalizable across different text-video encoders. To show this, we applied it to the Rank 1 encoder (Wang et al., 2025) from the Zero-Shot MSR-VTT leaderboard [^5] https://paperswithcode.com/sota/zero-shot-video-retrieval-on-msr-vtt as well as the MultiCLIP encoder (following Sanders et al. (2023)).
Table 1 shows the generality of our approach. When applied to the InternVideo2, our method improved the MSR-VTT NDCG score by 1 point without audio and 4 points with audio. Additionally, the significant performance drop observed for InternVideo2 on the MultiVENT dataset underscores the dataset’s richness and complexity. Despite this drop, incorporating our method boosted the NDCG score by 19 points without audio and 26 points with audio, further validating its effectiveness.
Model R@10 P@10 MRR NDCG
Arabic
MC 72.31 62.55 0.94 76.10
MC + Q2E 74.46 63.92 0.94 78.09
MC + Q2E + ASR 79.41 68.24 0.94 82.07
Chinese
MC 74.22 68.65 0.91 77.29
MC + Q2E 82.13 76.35 0.99 86.32
MC + Q2E + ASR 83.02 76.92 0.97 86.61
English
MC 85.79 81.92 1.00 89.98
MC + Q2E 86.45 82.50 1.00 90.57
MC + Q2E + ASR 87.80 83.85 1.00 91.43
Korean
MC 65.33 62.12 0.89 70.38
MC + Q2E 72.65 69.04 0.93 76.92
MC + Q2E + ASR 76.90 73.27 0.92 80.58
Russian
MC 79.09 71.15 0.96 82.84
MC + Q2E 81.61 73.46 0.96 84.62
MC + Q2E + ASR 85.40 76.54 0.99 88.70
Table 2: Comparison of different languages. Results are reported on the MultiVENT dataset. MC = MultiCLIP encoder. Results are reported on the MultiVENT dataset. The full table is reported on Table 9 in § A.1.4.
Conversely, MultiCLIP achieved the best baseline score on the MultiVENT dataset (Sanders et al., 2023). Applying our method to MultiCLIP resulted in an 8-point NDCG improvement, setting a new state-of-the-art for the MultiVENT dataset. However, unlike InternVideo2, MultiCLIP does not use a cross-encoder reranker, which explains its comparatively lower performance on the MSR-VTT dataset. Our method still led to a 4-point improvement in the MSR-VTT NDCG score, demonstrating effectiveness across different encoders.
Next, we examine the Q2E’s performance across languages (§ 5.3), the effect of different LLM sizes (§ 5.4), different rank fusion methods (§ 5.5) and various components of our method (§ 5.6). In the appendix, we examine the impact of VLM sizes (§ A.1.1), number of frames (§ A.1.2), and frame selection methods (§ A.1.3); Q2E is quite stable across encoding decisions.
5.3 Performance Across Languages
MultiVENT includes five languages – Arabic, Chinese, English, Korean, and Russian. Table 2 shows that building upon the MultiCLIP baseline, incorporating our method consistently improves retrieval performance across all metrics. Furthermore, adding audio features results in additional gains across all languages, demonstrating the effectiveness of our approach regardless of language.
MultiCLIP struggles the most with Arabic, Chinese, and Korean video retrieval. However, when our method is added, retrieval performance improves by approximately 6, 9, and 10 NDCG points for these languages, respectively.
5.4 Effect of LLM size
LLM Size R@10 P@10 MRR NDCG
MultiCLIP 70.82 65.25 0.92 75.34
Without Audio
1B 75.45 69.54 0.93 79.34
3B 75.50 69.54 0.94 79.78
8B 75.02 69.11 0.94 79.41
70B (Q2E) 75.76 69.73 0.95 80.04
With Audio
1B 78.71 72.28 0.95 82.50
3B 79.17 72.74 0.95 83.03
8B 79.04 72.59 0.95 82.91
70B (Q2E) 79.60 73.09 0.95 83.24
Table 3: Comparison of different LLM sizes within the Llama-3 family (1B to 70B parameters), specifically used for query decomposition and refining frame captions to video caption. Results are reported on the MultiVENT dataset. Full results in Table 10, § A.1.5.
Our primary model used the Llama-3.3-70B LLM. While this model is highly robust and generalizable, its large size demands significant computational power and increased inference time. To demonstrate the robustness of our method, Table 3 presents results for different LLM sizes, ranging from 1B to 70B parameters. Although the best performance is typically achieved with the 70B model, even the 1B model consistently outperforms the baseline NDCG by at least 4 NDCG points without audio and 8 points with audio. These results highlight the robustness of our method.
5.5 Effect of Rank Fusion Approaches
Method R@10 P@10 MRR NDCG
MultiCLIP 70.82 65.25 0.92 75.34
Without Audio
Neg. Exp. Ent. 60.05 55.60 0.90 66.61
RRF 62.91 58.30 0.92 69.74
Max 71.98 66.41 0.93 76.60
Mean 73.54 67.88 0.94 78.37
Inv. Ent. (Q2E) 75.76 69.73 0.95 80.04
With Audio
Neg. Exp. Ent. 67.23 61.97 0.93 73.20
RRF 70.91 65.29 0.93 76.29
Max 76.10 70.04 0.93 80.04
Mean 78.64 72.47 0.95 82.44
Inv. Ent. (Q2E) 79.60 73.09 0.95 83.24
Table 4: Comparison of aggregation methods for zero-shot rank fusion: Neg. Exp. Ent. (Negative Exponential Entropy), RRF (Reciprocal Rank Fusion), Inv. Ent. (Inverse Entropy), Mean and Max. Results are reported on the MultiVENT dataset. Full results in Table 11, § A.1.6.
Fusing different ranks in a zero-shot setting is a non-trivial task. Traditional approaches often involve training dense networks to combine scores (Yu et al., 2024) or leveraging LLMs to ensemble multiple expert outputs (Lu et al., 2024b, a; Jiang et al., 2023b). However, rank or score fusion is a significant challenge for LLMs, as they must process and compare a large volume of numerical data, including floating-point values (for scores) and discrete values (for ranks). In both cases, LLMs struggle to produce reasonable fusion ranks. We considered several aggregation methods, from simple techniques like Mean to more complex approaches such as Reciprocal Rank Fusion (RRF) and Inverse Entropy; due to space, we describe these formulations in § A.1.6. Results in Table 4 indicate that even the simple Mean function performs well compared to RRF and Negative Exponential Entropy. Inverse Entropy demonstrated better performance, leading us to adopt it as the aggregation method.
5.6 Effect of Different Components
Components R@10 P@10 MRR NDCG
MultiCLIP 70.82 65.25 0.92 75.34
Without Audio
Q2E 75.76 69.73 0.95 80.04
Q2E  – Video 58.00 53.90 0.89 64.83
Q2E  – Query 74.74 68.80 0.93 78.78
Q2E  – Events 74.77 68.88 0.93 79.02
With Audio
Q2E 79.60 73.09 0.95 83.24
Q2E  – Video 67.74 62.43 0.93 73.96
Q2E  – Query 78.12 71.74 0.93 81.54
Q2E  – Events 77.77 71.47 0.94 81.75
Table 5: Analysis of different components in our method. The first row in each block corresponds to our full method with all components included. In subsequent rows, one main component (video, query, event) is removed at a time to demonstrate its contribution to the overall method. Events = (Prequel + Current + Sequel). Results are reported on the MultiVENT dataset. The full result is reported on Table 12 in § A.1.7.
Table 5 presents an ablation study of query event decomposition and multimodal description components of Q2E. The results highlight that video is the most critical component. Removing the query results in a 1-point performance drop for both variants, while removing events reduces performance by 1 point in the non-audio variant and 2 points in the audio variant, highlighting the importance of decomposed events for retrieving relevant videos.
6 Conclusion & Future Work
We leveraged the world knowledge embedded in LLMs and VLMs to decompose both queries and videos for the video retrieval task, demonstrating how to match this knowledge against visual and audio signals. We employed inverse entropy as a rank fusion method without fine-tuning any additional fusion network. Experimental results across two datasets with different characteristics demonstrate the robustness and generalizability of our approach. While we show how the preexisting knowledge of LLMs and VLMs can be effectively transferred to text-to-video retrieval, our work sets the stage for future research to explore how factual and counterfactual information can be leveraged to positively or negatively align with the query.
7 Limitations
While our approach demonstrates superior performance across multiple datasets, it is computationally expensive and time-intensive. Future research should focus on optimizing efficiency to reduce both time and cost.
Due to the LLM and VLM’s inherent nature, some amount of misinformation or hallucination may be possible. Although we did not find any misinformation during our evaluation, our prompting of what could happen before, during, or after an event could result in hallucinations. Similarly, by relying on the parametric knowledge of these models, there is a potential for biases inherent in the models or their outputs to be propagated. We did not explicitly explore or mitigate the biases inherent in these models, though we also did not notice objectionable or stereotypical output.
8 Acknowledgement
This material is based in part upon work supported by the National Science Foundation under Grant No. IIS-2024878. Some experiments were conducted on the UMBC HPCF, supported by the National Science Foundation under Grant No. CNS-1920079. This material is also based on research that is in part supported by the Army Research Laboratory, Grant No. W911NF2120076, and by DARPA for the SciFy program under agreement number HR00112520301. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright notation thereon. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either express or implied, of ARL, DARPA or the U.S. Government.
Appendix A Appendix
A.1 Ablation Study (cont’d)
A.1.1 Effect of VLM Sizes
[…]
MultiCLIP 9.83 44.32 70.82 88.80 80.77 65.25 0.92 75.34 86.33 22.12 6
Without Audio
1B 9.80 43.76 71.73 89.19 79.85 66.02 0.93 75.86 86.05 24.96 6
2B 9.94 45.23 73.86 89.96 82.55 68.03 0.93 77.90 87.49 22.45 6
4B 9.97 46.75 75.85 90.73 85.17 69.92 0.93 79.81 89.05 19.68 6
8B 10.03 46.40 75.73 90.35 84.71 69.81 0.93 79.72 88.98 21.24 6
26B 10.24 47.03 76.16 92.66 85.64 70.19 0.95 80.47 89.94 19.87 6
38B (Q2E) 10.24 46.71 75.76 92.28 85.25 69.73 0.95 80.04 89.42 21.13 6
With Audio
1B 10.18 47.87 78.15 91.89 86.56 71.74 0.94 81.77 90.33 18.79 6
2B 10.14 48.27 78.46 91.51 87.26 72.05 0.94 82.14 90.58 17.21 6
4B 10.19 48.90 79.72 92.28 88.57 73.17 0.94 83.13 91.05 16.77 6
8B 10.40 48.87 79.88 93.82 88.34 73.36 0.95 83.52 91.53 17.27 6
26B 10.40 48.64 80.12 93.82 88.03 73.51 0.95 83.61 91.13 15.97 6
38B (Q2E) 10.32 49.00 79.60 93.05 88.73 73.09 0.95 83.24 91.20 16.44 6
Table 6: Comprehensive comparison of different VLM sizes within the InternVL-2.5 family (ranging from 1B to 38B parameters), specifically used for frame captioning. Results are reported on the MultiVENT dataset.
In Q2E, we utilized the InternVL-2.5-78B model. Similar to the Llama-3.3-70B, this model is highly robust and generalizable but comes with significant computational and time constraints. To assess the efficiency and flexibility of our method, Table 6 reports results across various VLM sizes within the InternVL-2.5 family.
The results indicate that models ranging from 4B to 38B parameters achieve nearly identical performance, with negligible differences in scores. This demonstrates that our method can achieve comparable results even with smaller VLMs, reducing computational requirements. Furthermore, with the inclusion of audio, the performance gap narrows even further; for instance, the 1B VLM experiences only a 1-point drop in NDCG score. These findings underscore the robustness and generalizability of our method, regardless of the VLM size.
A.1.2 Effect of Number of Frames
[…]
MultiCLIP 9.83 44.32 70.82 88.80 80.77 65.25 0.92 75.34 86.33 22.12 6
Without Audio
2 9.31 41.71 64.46 83.78 76.37 59.54 0.90 69.64 83.75 36.17 7
4 9.75 44.00 71.05 88.03 80.39 65.64 0.92 75.43 86.54 25.32 6
8 10.10 46.56 74.84 90.73 84.79 68.88 0.94 79.05 89.05 23.01 6
16 10.24 46.71 75.76 92.28 85.25 69.73 0.95 80.04 89.42 21.13 6
32 (Q2E) 10.15 47.22 75.95 91.89 85.95 70.00 0.94 80.29 90.15 20.05 6
64 10.09 47.15 76.77 91.12 85.95 70.66 0.94 80.66 89.74 19.98 6
With Audio
2 10.26 46.63 74.93 92.28 84.48 68.88 0.94 79.42 89.33 21.66 6
4 10.16 47.37 77.54 91.51 85.79 71.27 0.94 81.26 90.05 19.20 6
8 10.26 48.74 79.39 92.66 88.11 72.86 0.95 82.93 90.99 16.97 6
16 10.32 49.00 79.60 93.05 88.73 73.09 0.95 83.24 91.20 16.44 6
32 (Q2E) 10.41 48.83 80.09 93.82 88.26 73.55 0.95 83.68 91.32 16.38 6
64 10.40 48.90 80.42 93.82 88.49 73.82 0.95 83.95 91.58 15.90 6
Table 7: Full comparison of different number of frames (ranging from 2–64). Results are reported on the MultiVENT dataset.
Figure 3: Impact of the number of frames on retrieval performance. The figure shows the nDCG performance against comparisons per second for three methods: MultiCLIP (black), Q2E (w/o ASR) (blue), and the Q2E (w ASR) (orange). The point size is proportional to the number of sampled frames (i.e., 2, 4, 8, 16, 32, 64). Comparison per Second = Runtime / (num. of Query $\times$ num. of Video). Results are reported on the MultiVENT dataset. The full table is reported on Table 7.
Fig. 3 illustrates the impact of the number of frames on performance across different methods. The results show that increasing the number of sampled frames generally improves performance. However, as the number of comparisons per second increases (i.e., fewer frames are sampled), NDCG declines across all methods, highlighting a trade-off between computational efficiency and retrieval performance.
While the upward trend continues beyond 64 frames, we hypothesize that further increasing the number of frames would yield additional performance gains. However, the diminishing returns and increased computational cost would reduce the model’s practicality as a retrieval system. Therefore, we limited our evaluation to 64 frames and selected 16-frame sampling as the optimal balance between performance and speed.
A.1.3 Effect of Frame Selection Methods
[…]
MultiCLIP 9.83 44.32 70.82 88.80 80.77 65.25 0.92 75.34 86.33 22.12 6
Without Audio
scene_detect [^6] https://www.scenedetect.com/ 10.06 46.02 74.68 91.12 83.78 68.73 0.94 78.90 88.59 21.84 6
uniform (Q2E) 10.24 46.71 75.76 92.28 85.25 69.73 0.95 80.04 89.42 21.13 6
With Audio
scene_detect 10.36 48.55 79.34 93.44 87.95 72.78 0.95 82.99 90.98 17.58 6
uniform (Q2E) 10.32 49.00 79.60 93.05 88.73 73.09 0.95 83.24 91.20 16.44 6
Table 8: Comprehensive comparison of Different Frame Sampling Methods. In the scene-detect approach, we utilized the AdaptiveDetector with its default settings. For the uniform sampling method, we selected 16 evenly spaced frames, whereas the number of frames in the scene-detect approach varied based on the algorithm’s output. Results are reported on the MultiVENT dataset.
In Q2E, we employed a uniform sampling method for frame selection. However, to compare its effectiveness against a more complex approach, we conducted an ablation study using an adaptive detector from PySceneDetector [^7] https://www.scenedetect.com/. As shown in Table 8, uniform sampling achieves the best performance across both variants, demonstrating its robustness regardless of audio availability. Furthermore, the scene-detection algorithm yields a variable number of frames (ranging from 2 to 100), significantly increasing retrieval time by approximately 2.8 times compared to 16-uniform sampling.
A.1.4 Effect of Different Languages
[…]
Arabic
MultiCLIP 10.91 44.32 72.31 92.16 76.47 62.55 0.94 76.10 84.59 14.44 6.0
MultiCLIP + Q2E 10.82 45.86 74.46 90.20 78.43 63.92 0.94 78.09 86.08 14.91 6.0
MultiCLIP + Q2E + ASR 10.79 49.66 79.41 90.20 83.53 68.24 0.94 82.07 88.32 11.59 6.0
Chinese
MultiCLIP 9.12 45.59 74.22 84.62 83.85 68.65 0.91 77.29 86.52 12.60 6.0
MultiCLIP + Q2E 10.66 48.81 82.13 98.08 90.00 76.35 0.99 86.32 94.64 9.59 6.0
MultiCLIP + Q2E + ASR 10.46 50.55 83.02 96.15 93.08 76.92 0.97 86.61 93.96 8.73 5.5
English
MultiCLIP 10.61 50.82 85.79 100.00 96.15 81.92 1.00 89.98 97.47 7.64 5.0
MultiCLIP + Q2E 10.61 51.66 86.45 100.00 97.69 82.50 1.00 90.57 98.04 8.07 5.0
MultiCLIP + Q2E + ASR 10.61 51.66 87.80 100.00 97.69 83.85 1.00 91.43 98.12 7.74 5.0
Korean
MultiCLIP 9.30 39.98 65.33 86.54 75.00 62.12 0.89 70.38 82.86 20.78 7.0
MultiCLIP + Q2E 9.68 43.65 72.65 90.38 82.31 69.04 0.93 76.92 87.04 15.87 6.0
MultiCLIP + Q2E + ASR 9.68 45.93 76.90 90.38 86.15 73.27 0.92 80.58 89.02 13.85 6.0
Russian
MultiCLIP 10.57 49.57 79.09 92.31 87.69 71.15 0.96 82.84 91.20 10.65 6.0
MultiCLIP + Q2E 10.39 50.55 81.61 92.31 89.62 73.46 0.96 84.62 91.46 10.88 6.0
MultiCLIP + Q2E + ASR 11.21 53.49 85.40 98.08 94.23 76.54 0.99 88.70 95.02 9.13 5.0
Table 9: Comprehensive comparison of different languages. Results are reported on the MultiVENT dataset.
The detailed results for the experiment described in § 5.3 are reported on Table 9.
A.1.5 Effect of LLM Sizes
[…]
MultiCLIP 9.83 44.32 70.82 88.80 80.77 65.25 0.92 75.34 86.33 22.12 6
Without Audio
1B 10.05 46.07 75.45 90.35 83.94 69.54 0.93 79.34 88.15 22.34 6
3B 10.18 46.70 75.50 91.51 85.25 69.54 0.94 79.78 89.40 21.66 6
8B 10.21 46.60 75.02 92.28 84.94 69.11 0.94 79.41 89.37 19.74 6
70B (Q2E) 10.24 46.71 75.76 92.28 85.25 69.73 0.95 80.04 89.42 21.13 6
With Audio
1B 10.26 48.31 78.71 92.66 87.49 72.28 0.95 82.50 91.23 17.64 6
3B 10.32 48.88 79.17 93.05 88.57 72.74 0.95 83.03 91.55 16.51 6
8B 10.41 48.62 79.04 93.82 88.03 72.59 0.95 82.91 91.28 16.79 6
70B (Q2E) 10.32 49.00 79.60 93.05 88.73 73.09 0.95 83.24 91.20 16.44 6
Table 10: Comprehensive comparison of different LLM sizes within the Llama-3 family (ranging from 1B to 70B parameters), specifically used for query decomposition and refining frame captions to video caption. Results are reported on the MultiVENT dataset.
The detailed results for the experiment described in § 5.4 are reported on Table 10.
A.1.6 Effect of different Rank Fusion
[…]
MultiCLIP 9.83 44.32 70.82 88.80 80.77 65.25 0.92 75.34 86.33 22.12 6
Without Audio
Neg. Exp. Ent. 9.66 39.97 60.05 87.26 73.36 55.60 0.90 66.61 83.60 68.44 7
RRF 9.92 42.10 62.91 89.58 77.37 58.30 0.92 69.74 86.94 33.80 7
Max 9.98 45.15 71.98 90.35 82.32 66.41 0.93 76.60 87.77 29.31 6
Mean 10.19 45.99 73.54 92.28 84.09 67.88 0.94 78.37 89.50 26.32 6
Inv. Ent. (Q2E) 10.24 46.71 75.76 92.28 85.25 69.73 0.95 80.04 89.42 21.13 6
With Audio
Neg. Exp. Ent. 10.11 43.98 67.23 90.73 80.00 61.97 0.93 73.20 87.91 59.09 7
RRF 10.01 45.75 70.78 90.73 83.17 65.17 0.93 76.14 88.35 26.77 6
Max 9.94 46.58 76.20 89.96 84.79 70.12 0.93 79.86 88.82 22.89 6
Mean 10.23 48.06 78.30 92.66 87.41 72.16 0.95 82.20 91.04 20.32 6
Inv. Ent. (Q2E) 10.37 48.42 79.06 93.44 87.80 72.66 0.95 82.89 91.10 17.14 6
Table 11: Full comparison of different aggregation methods for zero-shot rank fusion: Neg. Exp. Ent. (Negative Exponential Entropy), RRF (Reciprocal Rank Fusion), Inv. Ent. (Inverse Entropy), Mean and Max. Results are reported on the MultiVENT dataset.
The detailed results for the experiment described in § 5.5 are reported on Table 11. These experiments consider five different approaches for aggregating the scores and performing zero-shot rank fusion. In all approaches, we compute a distribution $P_{i}$ across all videos, for each of the $i$ score components, as discussed in § 5.5. The specific rank fusion approaches we consider then vary in how we combine each of the five $P_{i}$ in order to get a single ranking $\hat{S}$ across all videos. Those approaches are:
Mean Aggregation
We simply average out $P_{i}$, computing $\hat{S}=\frac{1}{5}\sum_{i}P_{i}$.
Max Aggregation
For each video $v$, we maximize across the $P_{i}$ score for that video: $\hat{S}_{v}=\max_{i}P_{i,v}$.
Reciprocal Rank Fusion (RRF)
For each video $v$, we compute the rank of that video in $P_{i}$, and then sum across those reciprocal ranks. The final rank of $v$ is then $\sum_{i}(1/Rank(P_{i,v}))$.
Negative Exponential Entropy
We compute $\hat{S}=\sum_{i}\exp(-H(P_{i}))*P_{i}$, where $H(P_{i})$ is the entropy of $P_{i}$.
Inverse Entropy
This is the approach described in the main paper, in § 5.5.
A.1.7 Effect of Different Components
[…]
MultiCLIP 9.83 44.32 70.82 88.80 80.77 65.25 0.92 75.34 86.33 22.12 6
Without Audio
Q2E 10.24 46.71 75.76 92.28 85.25 69.73 0.95 80.04 89.42 21.13 6
Q2E  - Video 9.28 39.56 58.00 84.17 72.97 53.90 0.89 64.83 82.66 65.97 8
Q2E  - Query 10.08 45.97 74.74 91.12 83.78 68.80 0.93 78.78 87.92 21.00 6
Q2E  - Events 9.95 46.38 74.77 89.96 84.48 68.88 0.93 79.02 88.98 18.35 6
With Audio
Q2E 10.32 49.00 79.60 93.05 88.73 73.09 0.95 83.24 91.20 16.44 6
Q2E  - Video 10.16 44.61 67.74 91.51 81.16 62.43 0.93 73.96 88.97 50.96 6
Q2E  - Query 10.08 47.39 78.12 91.12 85.95 71.74 0.93 81.54 89.18 16.32 6
Q2E  - Events 10.22 48.09 77.77 91.89 87.18 71.47 0.94 81.75 90.38 16.64 6
Table 12: Full analysis of different components in our method. The top block shows the performance of our method with audio transcripts, while the bottom block represents the performance without audio transcripts. The first row in each block corresponds to our full method with all components included. In subsequent rows, one main component (query, video, event) is removed at a time to demonstrate its contribution to the overall method. Events = (Prequel + Current + Sequel). Results are reported on the MultiVENT dataset.
The detailed results for the experiment described in § 5.6 are reported on Table 12.
A.2 Dataset Statistics
The detailed dataset statistics are presented on Table 13.
Dataset # of query Avg. words # of video Avg. length (sec.)
MSR-VTT-1kA 995 9 1000 15
MultiVENT 259 27 2394 83
Table 13: Overview of the datasets used in this study. The table summarizes key statistics of the MSR-VTT-1kA and MultiVENT datasets.
A.3 Implementation Details
We used Llama-3888meta-llama/Llama-3.3-70B-Instruct as the LLM, InternVL-2.5999OpenGVLab/InternVL2_5-38B as the VLM, and Whisper-large-v3 101010openai/whisper-large-v3 as the multilingual ASR model for query decomposition, frame captioning and audio transcription. For embedding computations, we employed ColBERT111111hltcoe/plaidx-large for text-to-text similarity and MultiCLIP121212laion/Clip-VIT-H14 for text-to-video similarity score calculation. Query decomposition and video captioning were performed with a temperature of 0.8 and top-p sampling of 0.95. For frame extraction, we applied uniform sampling with 16 frames. Our evaluation framework is based on torchmetrics [^13] https://lightning.ai/docs/torchmetrics. The data extraction prompts were executed using two A100 GPUs, while the text-to-video retrieval was conducted on an RTX 6000 GPU.
A.4 Prompts
A.4.1 Extracting Spatial, Temporal, and Event Data
The following three prompts were utilized to extract the primary event, location, and time of the events from the query, respectively. The only input for these prompts is the original query. In cases where the query does not mention a specific time or location, the LLM was instructed to output “NOT AVAILABLE.”
[…]
[Extract Primary Event] You are given a video search query. Your task is to extract the primary event(s) being referenced or discussed in the query.
Here are some guidelines:
- Identify the main event(s) that the query is about, whether explicitly stated or implied
- If multiple events are mentioned, list all of them
- If no clear event can be identified from the query, output "NOT AVAILABLE" in the EVENTS section
- First, explain your reasoning in 1-2 sentences about how you identified the event(s) from the query
- Structure your response into two sections, "EXPLANATION:" and then "EVENTS:"
- The events section should contain either the identified event(s) in a numbered list, or "NOT AVAILABLE" if no event can be determined
- Don’t include temporal or spatial information in this event, just focus on the primary event(s) being discussed or referenced
Here is the query I’d like you to extract events from: {{query}}
[…]
[Extract Spatial Information] You are given a video search query. Assume the video search query is an ongoing event. Your task is to extract any locations mentioned in the query.
Here are some guidelines:
- Extract any location information mentioned in the query, including specific cities, towns, countries, regions, areas, neighborhoods, or other geographical references
- If no location information is found in the query, output "NOT AVAILABLE" in the LOCATION INFORMATION section
- First, explain your reasoning in 1-2 sentences about what location information appears in the query (or lack thereof) and why you’re extracting it
- Structure your response into two sections, "EXPLANATION:" and then "LOCATION INFORMATION:"
- The location information section should contain either the extracted location information in one line, or "NOT AVAILABLE" if no location information is provided.
- Output in the following format: CITY, COUNTRY, REGION, AREA, NEIGHBORHOOD, etc. (e.g. New York, USA, Europe, Downtown). If one of the fields is not available, you can leave it out.
Here is the query I’d like you to extract location information from: {{query}}
[…]
[Extract Temporal Informaton] You are given a video search query. Assume the video search query is an ongoing event. Your task is to extract any dates, times, or years mentioned in the query.
Here are some guidelines:
- Extract any temporal information mentioned in the query, including specific dates, times of day, years, seasons, or other time-related references
- If no temporal information is found in the query, output "NOT AVAILABLE" in the TEMPORAL INFORMATION section
- First, explain your reasoning in 1-2 sentences about what temporal information appears in the query (or lack thereof) and why you’re extracting it
- Structure your response into two sections, "EXPLANATION:" and then "TEMPORAL INFORMATION:"
- The temporal information section should contain either the extracted date/time information in one line, or "NOT AVAILABLE" if no temporal information is provided.
- Output in the following format: DAY MONTH YEAR, TIME, SEASON, etc. (e.g. 1 January 2022, 3:00 PM, Summer). If one of the fields is not available, you can leave it out.
Here is the query I’d like you to extract temporal information from: {{query}}
A.4.2 Decomposing Query to Prequel, Current, and Sequel Events
The following prompt was utilized to extract prequel, current, and sequel events from the query, with minimal modifications to maintain generalizability across all cases.
[…]
[ Prequel / Current / Sequel ] You are given a video search query. Assume the video search query is an ongoing event. You have to find out what previous events could lead to this given event / decompose the ongoing event in multiple simple events / find out what events can be the outcome of this ongoing event.The extracted events should be concrete enough to be visualized in a video.
Here are some guidelines:
- Extract events that could be visualized in a video.
- Before extracting an event, consider whether the event text could serve as a reasonable caption for a frame of a video associated with the query.
- Give specific events related to the query, use your previous WORLD KNOWLEDGE to extract events that are likely to be seen in a video. For example, if the query is "2019 Christmas Eve Protests in Hong Kong", you might return "Protesters marching in the streets", but not "Protestors marching in snow" because it is not likely to have snow in Hong Kong, though its pretty common in other places.
- First, explain your reasoning in 1-2 sentences about which events it makes sense to extract. Then, provide a numbered list with the events (anywhere from one to five event; use your best judgement to determine how many to extract)
- Structure your response into two sections, ""EXPLANATION:"" and then ""EVENTS:"". The events sections should contain nothing but the events.
Here is the query I’d like you to extract events from: {{query}}
A.4.3 Refining Query
After extracting information from the previous prompts, the following prompt was employed to refine the prequel, current, and sequel events by incorporating the extracted event, temporal, and spatial details. The goal was to construct natural, coherent, human-like search queries while preserving all relevant information.
[…]
[ Refine Query ] You will receive a query along with optional time, location, and main event information. Your task is to integrate this information naturally into a refined, human-like search query.
Here are some guidelines:
- Take the original query and incorporate any provided time, location, and event details seamlessly
- If any of these elements (time, location, event) are not provided, work with just the available information
- The refined query should sound natural, as if someone is searching with that query in Google, YouTube, etc.
- The base query information must always be preserved in the refined version
- First, explain your reasoning in 1-2 sentences about which events it makes sense to extract. Then, provide the refine query in one line
- Structure your response into two sections, ""EXPLANATION:"" and then ""REFINED QUERY:"". The Refined Query section should contain nothing but the refined query in one line.
Base Query: {{prequel/Current/sequel}}
Event: {{event}}
Place: {{place}}
Time: {{time}}
A.4.4 Frame Captioner
The following prompt was used to generate captions for each video frame. Unlike other prompts, a VLM was utilized to handle the multimodal inputs effectively. As illustrated in Fig. 2, we implemented two variants: (1) ASR and (2) Non-ASR. The Orange texts were excluded in the prompt for the Non-ASR variant.
[…]
[ Frame Caption ] You are provided with the description of the previous frames, the current frame (image) of a single video, and the original transcript of the audio (ASR) from the entire video. Your task is to describe the main event or activity depicted in the current frame.
# Instructions:
- Use the original ASR to enhance your understanding of the video context and inform your description of the current frame.
- Focus on identifying and detailing the key actions, participants, and context of the event depicted in the current frame.
- Maintain continuity with the previous frame’s description, connecting the current frame to the overall flow of the video.
- Ignore unrelated visual elements like colors or shapes unless they are critical to understanding the event.
- Use any visible text in the image to enhance your interpretation of the frame.
Only output the current frame description, integrating insights from the original ASR and previous frame description. Don’t output any additional commentary or content.
# Original ASR Transcript (Entire Video):
{{ASR}}
# Previous Frame Description:
{{prev_caption}}
# Frame:
{{frame}}
# Current Frame Description:
A.4.5 Full Video Captioner
The following prompt generated a summarized caption for the entire video. Given the high computational requirements and the limited capability of VLMs in handling videos, we utilized an LLM to summarize all the captions extracted in the previous step. The full refined audio transcript was also included in the ASR variant to provide additional context. As before, the Orange text elements were excluded from the prompt in the Non-ASR variant.
[…]
[ Video Caption using Frame Captions ] You will receive individual descriptions of consecutive frames from a single video, along with the original transcript of the audio (ASR) from the entire video. Your task is to provide a detailed description of the whole video, summarizing the key actions, participants, and context of the event depicted in the video.
# Instructions:
- Use the original ASR to enhance your understanding of the video context and inform your description of the whole video.
- Focus only on relevant aspects, such as what is happening, who is involved, and the overall context or setting.
- Ignore unrelated visual details like colors, shapes, or other non-essential information.
- Your summary should concisely and accurately capture the essence of the event, integrating insights from the frame descriptions, the Original ASR and the ASR summary.
Only output the video summary, without any additional commentary or content.
# Frame Descriptions:
## Frame [i] Description
{{frame_i_description}}
# Original ASR Transcript (Entire Video):
{{ASR}}
# Summary:
A.4.6 Refining Audio Transcripts
As illustrated in Fig. 2 (left orange box), we utilized an LLM to refine the translations generated by the multilingual-ASR and a translation-specific model. The following prompt was employed to refine the audio transcript using three inputs: (1) the original language transcript from ASR, (2) the translated transcript from the multilingual-ASR, and (3) the translated transcript from the translator.
[…]
[ Refine Audio Transcript ] You are an expert in language translation and text refinement. Your task is to produce a highly accurate and natural-sounding translation by analyzing the original transcript and two provided translations.
# Instructions:
- Combine the strengths of both translations and use the original transcript as a reference to improve accuracy and contextual meaning.
- Prioritize clarity, fluency, and maintaining the original intent and tone of the text.
- Where translations differ significantly, cross-check with the original transcript for fidelity.
- Ensure the final output is free from any awkward phrasing or mistranslation.
- If the transcripts can’t be refined with reasonable accuracy, just output "Not Available".
# Transcripts:
## Original Transcript:
{{original_language_asr}}
## Translation 1:
{{whisper_translated_asr}}
## Translation 2:
{{nllb_translated_asr}}
Output the refined translation only, without additional notes or explanations.
# Refined Translation:
A.5 Use of AI Assistance
The authors have used GitHub copilot  [^14] https://github.com/features/copilot during development and ChatGPT [^15] https://chatgpt.com/ for proofreading and polishing the final writing. Content provided to those tools were original to the authors.